# On an Erdős-Szekeres Game

### v2

@@ 7 Strategy for the achievement game
In most of this paper, we considered a permutation game where the first player to complete either an $I_{a}$ or a $J_{b}$ pattern loses. In this section we consider the same rules, except that the first player to complete either an $I_{a}$ pattern or a $J_{b}$ pattern wins.
@@
We describe the strategy in terms of the same boards as before. In both games, the first player who plays off of the $(b-1)\times(a-1)$ board ends the game. If this results in a loss, the goal is to play on the board as long as possible. The player who claims the lower right corner, indexed as $(a-1,b-1)$ is therefore the winner, since the entire board is eliminated, and their opponent must play off the board.
@@
However, if playing off the board results in a win, the penultimate move will be in column $a-1$ or row $b-1$, giving the player an opportunity to finish the appropriate monotone subsequence. Since playing in column $a-1$ or row $b-1$ gives the opponent a win, players seek to not play in this row and column. In other words, players wanting to win, try to restrict themselves to the $(b-2)\times(a-2)$ subboard. A player who claims cell $(a-2,b-2)$ then forces their opponent to play in the last row or last column, which allows the player who claimed $(a-2,b-2)$ to win. In other words, the achievement game requires the same strategy as the original game, but on a board with one fewer row and one fewer column. This means that player 1 has a winning strategy for the achievement game when $4\leq b\leq 6$. The parity of $a$ determines the winner when $b=3$. And player 2 is guaranteed a win for the achievement game when $b=2$ merely by completing a decreasing $J_{2}$ pattern on their first turn. This observation exactly matches Theorem 9 of [1].
8 Future Directions
In this paper we considered a particular two-player Erdős-Szekeres game. Unlike the game of Harary, Sagan, and West, where players chose fixed integers on each turn and where determining a strategy quickly became computationally untractable beyond a specific finite bound on $a+b$, in this game, players add any new $n$th digit to a permutation while preserving the pattern formed by the first $n-1$ digits. Playing this pattern-informed game allows for more generalized strategies where $a$ is arbitrarily large. Although the problem gets complex to analyze for sufficiently large $b$, we conjecture that a first player winning strategy exists for $a\geq b\geq 3$. We did, however, find a first player winning strategy for $a\geq b$ when $3\leq b\leq 5$.
In the remainder of this section we consider several directions for possible future work.
@@ 8.1 Permutations and boards
For sufficiently large $b$ it is convenient to phrase the strategies of this paper in terms of board shading rather than in terms of the actual underlying permutations in the original problem statement. However, at any point in the game, we may ask “how many permutations would have produced this particular board shading?” Clearly when $b=2$ and $b=3$, the permutations are unique. But for larger boards, multiple permutations would result in the same shaded board.
@@ 8.2 Number of moves
Table 1 shows the maximum and minimum number of moves possible in playing the Erdős-Szekeres game for $a\geq b$ and $2\leq b\leq 5$. The minimum is given by two players forming a decreasing permutation of length $b$. The maximum is given by the $(a-1)(b-1)+1$ bound given by Theorem 1. However the number of moves actually used by the strategies in this paper are in general somewhere between the two. When $b=2$, the two players create an $I_{a}$ permutation. When $b=3$, player 1 makes a strategic move to take cell $(a-1,2)$ which results in a final permutation one shorter than the maximum length in Theorem 1. When $b=4$, in general when player 2 makes a move that shades/eliminates 2 cells, player 1 responds by shading/eliminating 1 cell and vice versa. The difference between $12a-$ 1 or $32a-$ 3 total moves is decided by how player 2 handles the endgame. When $b=5$, although player 1 has a clear response to any move made by player 2, player 2 has far more options for what to do that have an impact on how quickly the game passes. At its slowest, players take turns moving/eliminating one cell at a time, and any turn that eliminates multiple cells come from the starting moves or the end game. At its fastest, player 2’s turn and player 1’s response shade or eliminate 4 cells, resulting in a permutation half as long as the maximum, minus a small finite number $C$ of extra cells eliminated by choices in the start and end game.
@@
A number of interesting follow up questions remain. For $b=5$, is there a more efficient winning strategy than the one presented here? For larger $b$, what strategies exist, and how do they compare, proportionally to the maximum length game guaranteed by Theorem 1?
@@ 8.3 Number of winning positions
Considering the board shading interpretation of the game results in additional interesting questions. We can shade a legal region of an $(a-1)\times(b-1)$ board in $\binom{(a-1)+(b-1)}{a-1}$ ways, by choosing the lattice path that divides the shaded and eliminated cells from the open cells. Table 2 gives data for various choices of $a$ and $b$. The first number is the number of shaded regions that result in a next-player loss, while the number in parentheses is this number divided by $\binom{(a-1)+(b-1)}{a-1}$; in other words, the second number is the percent of board shadings that result in a next-player loss (and thus are candidates to help form a set analogous to $\widehat{S}$ in the strategy for $a\geq b=5$. Of note, when $b$ is constant but $a$ increases, these percentages overall decrease. However, they show differences in parity, which is reflected in how the states of $\widehat{S}$ relied on how many open cells were in row 1. Do these values converge on a non-zero value as $a$ increases? If so, what is it?
$b\backslash a$ 2 3 4 5 6 7 8 9
@@
Table 2: Total number (and percentage) of shadings of an $(a-1)\times(b-1)$ board that result in a next-player loss

### v3

@@ 7.1 Strategy for normal play
In most of this paper, we considered a permutation game where the first player to complete either an $I_{a}$ or a $J_{b}$ pattern loses, i.e., the misère version of the game. Here we consider normal play, i.e., the first player to complete either an $I_{a}$ pattern or a $J_{b}$ pattern wins.
@@
We describe the strategy in terms of the same boards as before. In both games, the game is played on an $(b-1)\times(a-1)$ board. In misère play, the player who claims the lower right corner, indexed as (a1,b1)(a-1,b-1), is therefore the winner, since the entire board is eliminated, and their opponent must either complete an $I_{a}$ or $J_{b}$ pattern as the next move.
@@
However, in normal play, the penultimate move will be in column a1a-1 or row b1b-1, giving the player an opportunity to finish the appropriate monotone subsequence. Since playing in column a1a-1 or row b1b-1 gives the opponent a win, players seek to not play in this row and column. In other words, players wanting to win, try to restrict themselves to the $(b-2)\times(a-2)$ subboard. A player who claims cell (a2,b2)(a-2,b-2) then forces their opponent to play in the last row or last column, which allows the player who claimed (a2,b2)(a-2,b-2) to win. In other words, normal play requires the same strategy as misère play, but on a board with one fewer row and one fewer column. This means that player 1 has a winning strategy for normal play when $4\leq b\leq 6$. The parity of $a$ determines the winner when $b=3$. And player 2 is guaranteed a win for normal play when $b=2$ merely by completing a decreasing $J_{2}$ pattern on their first turn. This observation exactly matches Theorem 9 of [1].
More generally, Section 6 of [1] conjectures that normal play results in a first player win for $a\geq b\geq 4$. The analysis of this paper supports that conjecture by addressing the $b=6$ case directly. It remains to give a specific strategy for $b\geq 7$.
@@ 7.2 Permutations and boards
As $b$ increases, it is convenient to phrase the strategies of this paper in terms of board shading rather than in terms of the actual underlying permutations in the original problem statement. However, at any point in the game, we may ask “how many permutations would have produced this particular board shading?” Clearly when $b=2$ and $b=3$, the permutations are unique. But for larger boards, multiple permutations would result eliminating the same portion of the board.
@@ 7.3 Number of moves
Table 1 shows the maximum and minimum number of moves possible in playing the Erdős-Szekeres game for $a\geq b$ and $2\leq b\leq 5$. The minimum is given by two players forming a decreasing permutation of length $b$. The maximum is given by the (a1)(b1)+1(a-1)(b-1)+1 bound given by Theorem 1. However the number of moves actually used by the strategies in this paper are in general somewhere between the two. When $b=2$, the two players create an $I_{a}$ permutation. When $b=3$, player 1 makes a strategic move to take cell (a1,2)(a-1,2) which results in a final permutation one shorter than the maximum length in Theorem 1. When $b=4$, in general when player 2 makes a move that eliminates 2 cells, player 1 responds by eliminating 1 cell and vice versa. The difference between 2a12a-1 or 2a32a-3 total moves is decided by how player 2 handles the endgame. When $b=5$, although player 1 has a clear response to any move made by player 2, player 2 has far more options for what to do that have an impact on how quickly the game passes. At its slowest, players take turns eliminating one cell at a time, and any turn that eliminates multiple cells comes from the starting moves or the end game. At its fastest, player 2’s turn and player 1’s response eliminate 4 cells, resulting in a permutation half as long as the maximum, minus a small finite number $C$ of extra cells eliminated by choices in the start and end game.
@@
A number of interesting follow up questions remain. For $b=5$, is there a more efficient winning strategy than the one presented here? For larger $b$, what strategies exist, and how do they compare proportionally to the maximum length game guaranteed by Theorem 1?
@@ 7.4 Number of winning positions
Considering the board shading interpretation of the game results in additional interesting questions. We can shade a legal region of an $(a-1)\times(b-1)$ board in $\binom{(a-1)+(b-1)}{a-1}$ ways, by choosing the lattice path that divides the eliminated cells from the open cells. Table 2 gives data for various choices of $a$ and $b$. The first number is the number of game positions in class $\mathcal{P}$, while the number in parentheses is this number divided by $\binom{(a-1)+(b-1)}{a-1}$; in other words, the second number is the percent of game positions are in $\mathcal{P}$ (and thus are candidates to help form a set analogous to $\widehat{S}$ in the strategy for $a\geq b=5$). Of note, when $b$ is constant but $a$ increases, these percentages overall decrease. However, they show differences in parity, which is reflected in how the states of $\widehat{S}$ relied on the parity of open cells in row 1. Do these values converge on a non-zero value as $a$ increases? If so, what is it?
[…]
@@
Table 2: Total number (and percentage) of game positions on an $(a-1)\times(b-1)$ board in class $\mathcal{P}$
Acknowledgements
The author is grateful to two anonymous referees who provided detailed and helpful feedback that improved the presentation of this paper.